OpenStudy (anonymous):

What is the length of side AB in the triangle below? 19.21 9 3.87 4.12

OpenStudy (anonymous):

4,12

OpenStudy (anonymous):

sqrt((-8--4)^2 + (-8-7)^2) = sqrt((-4)^2 + (-1)^2) = sqrt 17 = 4.12